OpenStudy (anonymous):

Assume the following respective probabilities for the positions of the two loads: P(W1 at B) : 0.25 P(W1 at C) : 0.60 P(W2 at B) : 0.30 P(W2 at C): 0.50 Assuming that the positions of W1 and W2 are statistically independent; what are the respective probabilities associated with each of the possible values of MA.

OpenStudy (kropot72):

What does "MA" signify?

OpenStudy (anonymous):

this is the whole excercise:

OpenStudy (kropot72):

Looks as if Ma is the value of the bending moment. Now it makes sense :)

OpenStudy (kropot72):

An important property of independent events is that the probability of both occurring is the product of their probabilities. The formula is: \[P(A \cap B)=P(A)\times P(B)\] Therefore we find that \[P(W1+W2\ at\ B)=0.25\times 0.30=?\] Do you follow this?

OpenStudy (kropot72):

If W1 and W2 are both at point B, the total load is 200 + 500 = 700 lb at a distance of 10 ft from the mounting. Therefore the bending moment is \[700\times 10=7000\ lb-ft\] The probability of this bending moment is \[P(M _{A}=7000)=0.25\times 0.30=you\ can\ calculate\]

OpenStudy (kropot72):

@marleisa.arocho Are you there?

OpenStudy (anonymous):

and if there is only \[W _{1} at B\]

OpenStudy (kropot72):

It is given that \[P(W _{1}\ at\ B)=0.25\] The bending moment is \[200\times 10=2000\ lb-ft\] \[P(M _{A}=2000)=0.25\]